OpenStudy (anonymous):

Solve for the variable. -a/4=-15

OpenStudy (anonymous):

Multiply both the sides by -4 and tell me what did you get??

OpenStudy (anonymous):

Like:\[\frac{-a}{4} \times (-4) = -15 \times (-4)\]

OpenStudy (anonymous):

@heartbreak tell me what will you get when you will do the above..

OpenStudy (anonymous):

-60

OpenStudy (anonymous):

-(-) = +.. Firstly solve LHS..

mathslover (mathslover):

\[\large{-a\times -b=ab}\]

mathslover (mathslover):

@heartbreak Always remember that :when two negatives are multiplied then the result is Positive : \[\large{-2\times -3=6 \ne -6}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what did you get now ? @heartbreak

mathslover (mathslover):

so can you tell me again : \(\large{-15\times -4=?}\)

OpenStudy (anonymous):

+60

mathslover (mathslover):

very great..

OpenStudy (anonymous):

Yes.. Well done..

mathslover (mathslover):

and that is your answer..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear..

mathslover (mathslover):

also : \[\LARGE{\mathbb{Welcome}\textbf{To}\mathbb{Open}\textbf{Study}}\] Best of luck for your journey in openstudy